OpenStudy (anonymous):

The length of Johns patio is 5 feet less than 3 times the length of Howards patio. Johns patio has a length of 40 feet. What ias the length, l, of Howards patio?

OpenStudy (texaschic101):

J = 3H - 5 J = 40 now sub 40 in for J 40 = 3H - 5 40 + 5 = 3H 45 = 3H 15 = H check... 40 = 3(15) - 5 40 = 45 - 5 40 = 40 (correct) Howard's patio is 15 ft long